# Bridgeland/weak Stability Conditions under Spherical Twist Associated to a Torsion Sheaf

### v1

@@ 6. Stability of some line bundles at $\sigma^{b}_{V,\nu}$
In this Section, we study the stability of certain line bundles at $\sigma^{b}_{V,\nu}$. Let $X$, $\nu$, $C$ be as in Section 4. We begin with an observation of image of line bundles under spherical twist. Proposition 6.1. Let $L$ be an line bundle on $X$, and $c:=\mathrm{ch}_{1}(L)\cdot C$. Then we have
(i) If $t-c=-1$, then $\text{ST}_{\mathcal{O}_{C}(t)}(L)=L$.
(ii) If $t-c<-1$, then $\text{ST}_{\mathcal{O}_{C}(t)}(L)$ is a coherent sheaf of rank $1$. In particular, if $t-c=-2$, $\text{ST}_{\mathcal{O}_{C}(t)}(L)=L(C)$.
(iii) If $t-c<-1$, then $\text{ST}_{\mathcal{O}_{C}(t)}^{-1}(L)$ is a coherent sheaf of rank $1$.
Proof.
By definition, $\text{ST}_{\mathcal{O}_{C}(t)}(E)$ fits in the following exact triangle:
$\textnormal{RHom}(\mathcal{O}_{C}(t),E)\otimes^{L}\mathcal{O}_{C}(t)\rightarrow E\rightarrow\text{ST}_{\mathcal{O}_{C}(t)}(E)\xrightarrow{+1}.$
We have
$\begin{split}\textnormal{Ext}^{i}(\mathcal{O}_{C}(t),L)&\simeq\textnormal{Ext}^{2-i}(L,\mathcal{O}_{C}(t))^{*}\\ &\simeq H^{2-i}(\mathcal{O}_{C}(t-c))^{*}.\end{split}$
When $t-c=-1$, $H^{2-i}(\mathcal{O}_{C}(t-c))^{*}=0$ for all $i$, hence $\text{ST}_{\mathcal{O}_{C}(t)}(L)=L$.
When $t-c<-1$, the only nonzero cohomology is given by $H^{1}(\mathcal{O}_{C}(t-c))\simeq\textnormal{Ext}^{1}(\mathcal{O}_{C}(t),L)$. Hence we have
$0\rightarrow L\rightarrow\text{ST}_{\mathcal{O}_{C}(t)}(L)\rightarrow\mathcal{O}_{C}(t)^{n}\rightarrow 0,$
where $n=\text{dim }\textnormal{Ext}^{1}(\mathcal{O}_{C}(t),L)$. Hence $\text{ST}_{\mathcal{O}_{C}(t)}(L)$ is a coherent sheaf of rank $1$. When $t-c=-2$, we have $n=1$, which implies that $\text{ST}_{\mathcal{O}_{C}(t)}(L)\simeq L(C)$.
Similarly, the formular of $\text{ST}_{\mathcal{O}_{C}(t)}^{-1}$ implies that
$\text{ST}_{\mathcal{O}_{C}(t)}^{-1}(L)\to L\to\mathcal{O}_{C}(t)[2]\otimes^{L}\textnormal{RHom}(\mathcal{O}_{C}(t),L)\xrightarrow{+1}.$
When $t-c\leq-1$, $\text{ST}_{\mathcal{O}_{C}(t)}^{-1}(L)$ fits into the following s.e.s in $\mathrm{Coh}(X)$:
(6.1.1) $0\to\mathcal{O}_{C}(t)^{n}\to\text{ST}_{\mathcal{O}_{C}(t)}^{-1}(L)\to L\to 0$
@@
(6.2.2) $0\to K^{\prime}\to A\to F^{\prime}\to 0$
Then $F^{\prime}\in\mathcal{T}^{0}_{\nu}$, and $\textnormal{Hom}(A,\mathcal{C}_{-2})=0$ implies that $\textnormal{Hom}(F^{\prime},\mathcal{C}_{-2})=0$. Hence $F^{\prime}\in\mathcal{T}_{\nu,-2}$. Since $K^{\prime}\in\mathcal{T}_{\nu,-2}$, we have 6.2.2 is a s.e.s in $\mathcal{B}_{\nu,-2}$.
@@
Combining Equation 6.2.3 and 6.2.4, we have
@@
Since $\mu_{\nu}(A)\geq\mu_{\nu}(F)$, we have
(6.2.5) $\mathrm{ch}_{0}(A)\leq\frac{\alpha\cdot\nu}{\nu^{2}(\sqrt{(\frac{1/2\alpha^{2}-V}{\alpha\cdot\nu})^{2}+\frac{2}{\nu^{2}}V}+\frac{1/2\alpha^{2}-V}{\alpha\cdot\nu})}$
@@
in $\mathcal{B}_{\nu,-2}$ such that $A$ satisfies (i’) $\phi_{V_{0},\nu}(A)=\phi_{V_{0},\nu}(L)=\phi_{V_{0},\nu}(Q)$ and (ii) in Lemma 6. Proposition 6.4. Assume that $\nu\in\text{NS}_{\mathbb{Q}}(X)$, Then the set of destabilizing walls $W(A,L)$ is locally finite.
@@
The argument follows from Proposition 3.10 in [10]. Let $I=[a,b]$ be an interval in the ray $V\cdot\nu$. Then a numerical wall $W(A,L)$ intersect $I$ at points
$V=\frac{\mathrm{ch}_{2}(A)(\alpha\cdot\nu)-\frac{1}{2}\alpha^{2}(\mathrm{ch}_{1}(A)\cdot\nu)}{\mathrm{ch}_{0}(A)(\alpha\cdot\nu)-\mathrm{ch}_{1}(A)\cdot\nu}.$
Since $\nu\in\text{NS}_{\mathbb{Q}}(X)$, we know that $V\in I\cap\mathbb{Q}$. To show that there are only finitely many possibilities for $V$ in $I\cap\mathbb{Q}$, we only need to show that the term $|\mathrm{ch}_{0}(A)(\alpha\cdot\nu)-\mathrm{ch}_{1}(A)\cdot\nu|$ is bounded above.
By Lemma 6, we know that there exists $N_{I}\in\mathbb{Z}$ such that $0<\mathrm{ch}_{0}(A)<N_{I}$. Since $0<\mathrm{ch}_{1}(A)\cdot\nu\leq\alpha\cdot\nu$, the term $|\mathrm{ch}_{0}(A)(\alpha\cdot\nu)-\mathrm{ch}_{1}(A)\cdot\nu|$ is bounded above.
By the previous two lemmas, the outermost destabilizing wall for higher rank object exists, and we denote it by $V_{h}$. Assume that there exist $0\to G\to L\to Q\to 0$ in $\mathcal{B}_{\nu,-2}$ such that $\phi^{b}_{V_{1},\nu}(G)=\phi^{b}_{V_{1},\nu}(L)=\phi^{b}_{V_{1},\nu}(Q)$, $\mathrm{ch}_{0}(G)>1$ and $V_{1}>V_{h}$. Then $G$ is not $\sigma^{b}_{V_{1},\nu}$-semistable. If $G$ has a torsion subsheaf, we can quotient it out and obtain a torsion free sheaf with a higher wall, hence we may assume that $G$ is torsion free. Consider the s.e.s in $B_{\nu,-2}$
$0\to G_{1}\to G\to G^{\prime}\to 0$
where $G_{1}=\text{HN}_{1}(G)$ at $\sigma^{b}_{V_{1},\nu}$. Then we have $\phi^{b}_{V_{1},\nu}(G_{1})\geq\phi^{b}_{V_{1},\nu}(G)$, which forces $\mathrm{ch}_{0}(G_{1})=1$. Hence in this situation, $\phi^{b}_{V_{1},\nu}(G_{1})\geq\phi^{b}_{V_{1},\nu}(L)$ and $W(G,L)$ cannot be the outermost wall for $L$. Hence for a higher rank object $G\hookrightarrow L\in\mathcal{B}_{\nu,-2}$, if $W(G,L)$ is the outermost wall, then $V_{1}=W(G,L)\leq V_{h}$.
If there exist $V^{b}_{u}$ such that $L$ is strictly semistable at $\sigma^{b}_{V^{b}_{u},\nu,uC}$, and $L$ is unstable at $\sigma^{b}_{V,\nu,uC}$ when $V<V^{b}_{u}$, we say the outermost wall along $\sigma^{b}_{V,\nu,uC}$ exists and equals to $V^{b}_{u}$.
Proposition 6.5. Let $L$ be a line bundle on $X$ with $c=-1$ or 00. Let $10<u<$ 1 and assume the outermost wall along $\sigma^{b}_{V,\nu,uC}$ exists. Assume that $0\to E\to L\to Q\to 0$ is a s.e.s. in $\mathcal{B}_{\nu,-2}$ such that $W(E,L)$ along $\sigma^{b}_{V,\nu,uC}$ equals to $V^{b}_{u}$. Let $n_{i}(E):=\textnormal{Ext}^{i}(\mathcal{O}_{C}(-1),E)$, then $n_{0}(E)=n_{1}(E)=0$, $n_{2}(E)=0$ or $1$.
@@
First note that $L\in\mathcal{B}_{\nu,-2}$ implies that $c>-2$. Let $E\hookrightarrow L\in\mathcal{B}_{\nu,-2}$ such that $W(E,L)$ along $\sigma^{b}_{V,\nu,uC}$ equals to $V^{b}_{u}$. By the discussion of the paragraph above, $E$ must be torsion free, which implies that $n_{0}(E)=0$.
Next we check that $n_{1}(E)=0$. Assume not, we have an extension
$0\to E\to E^{\prime}\to\mathcal{O}_{C}(-1)\to 0$
in $\mathcal{B}_{\nu,-2}$, the assumption on $c$ implies that the map $E\to L$ induces a map $E^{\prime}\to L$. If $\ker(E^{\prime}\to L)\neq 0$, then snake lemma and $\textnormal{Hom}(\ker(E^{\prime}\to L),\mathcal{O}_{C}(-2))=0$ implies that $\ker(E^{\prime}\to L)\simeq\mathcal{O}_{C}(-1)$. Since
$\textnormal{Hom}(\mathcal{O}_{C}(-1),\mathcal{O}_{C}(-1))\to\textnormal{Ext}^{1}(\mathcal{O}_{C}(-1),E)$
is nonzero, we have $\textnormal{Hom}(\mathcal{O}_{C}(-1),E^{\prime})\to\textnormal{Hom}(\mathcal{O}_{C}(-1),\mathcal{O}_{C}(-1))$ is the zero map. This forces $\ker(E^{\prime}\to L)=0$ and $E^{\prime}$ is also a subobject of $L$ in $\mathcal{B}_{\nu,-2}$. Furthermore, $u>0$ implies that $\phi^{b}_{V,\nu,uC}(E^{\prime})>\phi^{b}_{V,\nu,uC}(E)$, hence $W(E^{\prime},L)$ is higher than $V^{b}_{u}$, contradicting $V^{b}_{u}$ is the outermost wall.
Denote the image of $E\to L$ in $\mathrm{Coh}(X)$ by $L_{n}(-C^{\prime})$. Consider the s.e.s in $\mathcal{A}_{\nu,-2}$:
$0\to\text{ST}(E)\to\text{ST}(L)\to\text{ST}(Q)\to 0$
By the l.e.s. of cohomology, we have $H^{0}(\text{ST}(E))\simeq E$ and $H^{0}(\text{ST}(L))\simeq L$. This implies that $H^{-1}(\text{ST}(Q))\simeq H^{-1}(Q)$.
If $c=0$, we also have a s.e.s. in $\mathrm{Coh}(X)$:
$0\to H^{0}(Q)\to H^{0}(\text{ST}(Q))\to\mathcal{O}_{C}(-1)^{n_{2}}\to 0$
Since $\textnormal{Hom}(L_{n}(-C^{\prime}),\mathcal{O}_{C}(-2))=0$, $\dim\textnormal{Hom}(L_{n}(-C^{\prime}),\mathcal{O}_{C}(-1))\leq 1$. Applying $\textnormal{Hom}(\mathcal{O}_{C}(-1),\_)$ to
$0\to L_{n}(-C^{\prime})\to L\to H^{0}(Q)\to 0,$
we obtain $\dim\textnormal{Ext}^{1}(H^{0}(Q),\mathcal{O}_{C}(-1))\leq 1$. Hence $n_{2}\leq 1$.
If $c=-1$, we have a s.e.s. in $\mathrm{Coh}(X)$:
$0\to H^{0}(Q)\to H^{0}(\text{ST}(Q))\to\mathcal{O}_{C}(-1)^{m}\to 0$
where $m=n_{2}$ or $n_{2}-1$. If $\textnormal{Hom}(Q,\mathcal{O}_{C}(-1))=0$, we have $\textnormal{Hom}(H^{0}(Q),\mathcal{O}_{C}(-1))=0$ and $m=n_{2}-1$. In this case, we have $\textnormal{Ext}^{1}(H^{0}(Q),\mathcal{O}_{C}(-1))=0$, hence $m=0$ and $n_{2}=1$. If $\textnormal{Hom}(Q,\mathcal{O}_{C}(-1))\simeq\mathbb{C}$, then $m=n_{2}$ and $\dim\textnormal{Ext}^{1}(H^{0}(Q),\mathcal{O}_{C}(-1))\leq 1$, hence $m\leq 1$ and $n_{2}\leq 1$.
Remark 6.6. If the outermost wall along $\sigma^{b}_{V,\nu}$ exists, then there exists $E\hookrightarrow L\in\mathcal{B}_{\nu,-2}$ such that $W(E,L)=V_{0}^{b}$ and $E$ satisfies the same conclusion in Proposition 6. By the same reasoning as in Proposition 6, for any $E\hookrightarrow L\in\mathcal{B}_{\nu,-2}$ such that $W(E,L)=V_{0}^{b}$, we have $n_{0}(E)=0$ and $n_{2}\leq 1$. Let $E\hookrightarrow L\in\mathcal{B}_{\nu,-2}$, and assume that $n_{1}(E)\neq 0$ Then we form
$0\to E\to E^{\prime}\to\mathcal{O}_{C}(-1)^{n_{1}(E)}\to 0$
Then $E^{\prime}$ is also a subobject of $L$ in $\mathcal{B}_{\nu,-2}$, $n_{1}(E^{\prime})=0$ and $W(E^{\prime},L)=V^{b}_{0}$.
Proposition 6.7. Let $L$ be a line bundle on $X$ such that $c=0$. Assume the outermost wall along $\sigma^{b}_{V,\nu}$ exists. Let $0\to E\to L\to Q\to 0$ be a s.e.s. in $\mathcal{B}_{\nu,-2}$ such that $W(E,L)$ along $\sigma^{b}_{V,\nu}$ equals to $V^{b}_{0}$. Then we have $C\cdot\mathrm{ch}_{1}(H^{0}(Q))\geq-2$
@@
Since $H^{0}(Q)$ is a torsion sheaf, we can write $\mathrm{ch}_{1}(H^{0}(Q))=a_{0}C+\sum a_{i}C_{i}$ where $a_{i}\in\mathbb{Z}_{\geq 0}$ and $C_{i}$’s are irreducible curves in $X$ which does not equal to $C$. If $a_{0}=0$, then the conclusion is automatically true. From now on we assume $a_{0}\neq 0$. Note that $H^{0}(Q)\otimes L^{*}$ is the structure sheaf of a subscheme of $X$, consider the connected component of $C$ in the subscheme, and denote it by $Y$. Then we have $\mathcal{O}_{Y}|_{C}\simeq\mathcal{O}_{C}$, denote $K_{1}:=\ker(\mathcal{O}_{Y}\to\mathcal{O}_{C})$. Then $\mathcal{O}(-C)\twoheadrightarrow K_{1}$, which implies that $\mathcal{O}_{C}(2)\twoheadrightarrow K_{1}$, then if nonzero, $K_{1}|_{C}\simeq\mathcal{O}_{C}(2)$ or is supported on dimension 00. If $K_{1}|_{C}\simeq\mathcal{O}_{C}(2)$, denote $\ker(K_{1}\to\mathcal{O}_{C}(2))$ by $K_{2}$ and repeat this process, we get $K_{a_{0}-1}|_{C}\simeq\mathcal{O}_{C}(2(a_{0}-1))$, and $K_{a_{0}}=0$ or supported in dimension 00. Then we have $H^{0}(Q)$ fits into a s.e.s.
$0\to M\otimes L\to H^{0}(Q)\to N\to 0,$
where $M|_{C}$ is zero or supported in dimension zero, $N\in\langle\mathcal{O}_{C}(i)|i\geq 0\rangle$. Then
$\textnormal{Hom}(M\otimes L,\mathcal{O}_{C}(-1))=\textnormal{Hom}(M,\mathcal{O}_{C}(-1))=0.$
Using local to Global spectral sequence, for $i\geq 1$ we have $\textnormal{Ext}^{1}(\mathcal{O}_{C}(i),\mathcal{O}_{C}(-1))\simeq\mathbb{C}^{i}$. Consider the s.e.s.
$0\to\mathcal{O}_{C}(2)\to P\to\mathcal{O}_{C}\to 0$
Since $\dim\textnormal{Hom}(\mathcal{O}_{C}(-1),P)=6$, we have $\textnormal{Ext}^{1}(P,\mathcal{O}_{C}(-1))\neq 0$. Hence if $a_{0}>1$, $\dim\textnormal{Ext}^{1}(N,\mathcal{O}_{C}(-1))\geq 2$. This implies that $\dim\textnormal{Ext}^{1}(H^{0}(Q),\mathcal{O}_{C}(-1))\geq 2$. But from the proof of the previous Proposition, we have $\dim\textnormal{Hom}(L_{n}(-C^{\prime}),\mathcal{O}_{C}(-1))\leq 1$. Then the l.e.s. of cohomology implies that
$\dim\textnormal{Ext}^{1}(H^{0}(Q),\mathcal{O}_{C}(-1))\leq 1,$
and we obtain a contradiction. Hence $a_{0}\leq 1$.
Proposition 6.8. Let $X$ be a smooth K3 surfaces. Let $L$ be a line bundle on $X$ such that $c=0$. Assume that for any irreducible negative curve $C^{\prime}\subseteq X$, $|C^{\prime}\cdot C|\geq 2$. Then the outermost wall along $\sigma^{b}_{V,\nu}$ does not exist, and one of the following cases holds:
(i) $L$ is $\sigma^{b}_{V,\nu}$-stable for all $V>0$.
(ii) $L$ is unstable at $\sigma^{b}_{V,\nu}$ for some $V>0$. Then for any $E\hookrightarrow L\in\mathcal{B}_{\nu,-2}$, there exists $F\hookrightarrow L\in\mathcal{B}_{\nu,-2}$ and $\mathrm{ch}_{0}(F)=1$, such that $W(E,L)$ is inside $W(F,L)$.
Proof.
Assume the outermost wall exists. Let $E\hookrightarrow L\in\mathcal{B}_{\nu,-2}$ such that $W(E,L)=V^{b}_{0}$. Write $\mathrm{ch}_{1}(L)=\alpha\nu+\beta H$, where $H\cdot\nu=0$. Consider the space of Bridgeland/weak stability conditions $\sigma^{b}_{\nu,B}$, where $B=y\nu+zH$. By the same argument in [1], we have $C^{\prime 2}<0$. Comparing the centers of the numerical walls, we have when $|z|\gg 0$, $W(E,L)$ is lower than $W(L(-C^{\prime\prime}),L)$ for some irreducible curve $C^{\prime\prime}$ with $C^{\prime\prime 2}<0$.
@@
Since $c=0$, we have $H\cdot C=0$, and hence $B\cdot C=0$. Then $\phi^{b}_{\nu,B}(L_{n}(-C))<\phi^{b}_{\nu,B}(L)$ for any $B=y\nu+zH$. Then we have $C^{\prime\prime}\neq C$, and $C^{\prime\prime}\cdot C\geq 0$. On the other hand, $\dim\textnormal{Hom}(L_{n}(-C^{\prime}),\mathcal{O}_{C}(-1))\leq 1$ implies that $\dim\textnormal{Hom}(L(-C^{\prime}),\mathcal{O}_{C}(-1))\leq 1$ and $\dim\textnormal{Hom}(L(-C^{\prime\prime}),\mathcal{O}_{C}(-1))\leq 1$, hence $C^{\prime\prime}\cdot C\leq 1$. Combine these two results, we have $|C^{\prime\prime}\cdot C|\leq 1$, by assumption of the Proposition, such $C^{\prime\prime}$ does not exists.
@@
The K3 surfaces satisfies the assumption in the Proposition 6 exists. Example 6.9. Consider the case Theorem 2(v) in [9]. Let $\rho(X)=2$. Then the Effective cone of $X$ is generated by two smooth rational $(-2)$ curves. Denoting the two curves by $C_{1}$ and $C_{2}$ and $C_{1}\cdot C_{2}=q$, then $q>2$. Since otherwise, the first Chern character of any line bundle has non-positive self-intersection, which implies that there is no ample line bundle on $X$, a contradiction. In this case, take $y$ to be an even positive integer, and $\nu=\frac{q}{2}yC_{1}+yC_{2}$, then $\nu$ satisfies $\nu\cdot C_{1}=0$, $\nu\cdot C_{2}>0$ and $\nu^{2}>0$. Example 6.10. Again consider the case Theorem 2(v) in [9]. Let $\rho(X)=3$. Assume that the effective cone of $X$ is generated by three curves $C_{1}$, $C_{2}$ and $C_{3}$. Then Corollary 2.9 (i) in [11] implies that there exists K3 surface such that the pairwise intersection between $C_{i}$’s are greater than $2$.

### v2

@@ 6. Stability of some line bundles at $\sigma^{b}_{V,\nu}$
In this Section, we study the stability of certain line bundles at $\sigma^{b}_{V,\nu}$. Let $X$, $\nu$, $C$ be as in Section 4.
@@
(6.1.2) $0\to K^{\prime}\to A\to F^{\prime}\to 0.$
Then $F^{\prime}\in\mathcal{T}^{0}_{\nu}$, and $\textnormal{Hom}(A,\mathcal{C}_{-2})=0$ implies that $\textnormal{Hom}(F^{\prime},\mathcal{C}_{-2})=0$. Hence $F^{\prime}\in\mathcal{T}_{\nu,-2}$. Since $K^{\prime}\in\mathcal{T}_{\nu,-2}$, we have 6.1.2 is a s.e.s in $\mathcal{B}_{\nu,-2}$.
@@
Combining Equation 6.1.3 and 6.1.4, we have
@@
Let $\alpha:=\mathrm{ch}_{1}(L)$. Since $\mu_{\nu}(A)\geq\mu_{\nu}(F)$, we have
(6.1.5) $\mathrm{ch}_{0}(A)\leq\frac{\alpha\cdot\nu}{\nu^{2}(\sqrt{(\frac{\frac{1}{2}\alpha^{2}-V}{\alpha\cdot\nu})^{2}+\frac{2}{\nu^{2}}V}+\frac{\frac{1}{2}\alpha^{2}-V}{\alpha\cdot\nu})}$
@@
in $\mathcal{B}_{\nu,-2}$ such that $A$ satisfies (i’) $\phi_{V_{0},\nu}(A)=\phi_{V_{0},\nu}(L)=\phi_{V_{0},\nu}(Q)$ and (ii) in Lemma 6. Using the same argument of Proposition 3.10 in [12], we have the set of destabilizing walls $W(A,L)$ is locally finite for $\nu\in\text{NS}_{\mathbb{Q}}(X)$. By the previous lemma, the highest destabilizing wall for higher rank object exists, and we denote it by $V_{h}$.
Proposition 6.3. Let $L$ be a line bundle on $X$, then for each $E\hookrightarrow L$ in $\mathcal{B}_{\nu,-2}$, there exists $E^{\prime}\hookrightarrow L$ in $\mathcal{B}_{\nu,-2}$ such that $\textnormal{Hom}(E^{\prime},\mathcal{O}_{C}(-1))=0$ and $\sigma^{b}_{V,\nu}(E)=\sigma^{b}_{V,\nu}(E^{\prime})$.
@@
Since $E\in\mathcal{B}_{\nu,-2}$, $\textnormal{Hom}(E,\mathcal{O}_{C}(-2))=0$. If $\textnormal{Hom}(E,\mathcal{O}_{C}(-1))\neq 0$, any map from $E$ to $\mathcal{O}_{C}(-1)$ must be surjective in $\mathrm{Coh}(X)$. We complete to a s.e.s. in $\mathrm{Coh}(X)$:
(6.3.1) $0\to E_{1}\to E\to\mathcal{O}_{C}(-1)\to 0.$
Consider the following commutative diagram formed by pullout:
[…]
where $\text{HN}_{n}(E_{1})$ is the last HN factor of $E_{1}$, $P$ is formed by pushout diagram, then $R$ is a quotient sheaf of $\mathcal{O}_{C}(-1)$. Since $\mathcal{T}_{\nu,-2}$ is a torsion class in $\mathrm{Coh}(X)$, we have $P\in\mathcal{T}_{\nu,-2}$. Then $\mu_{\nu}(R)=0$ implies that $\mu_{\nu}(\text{HN}_{n}(E_{1}))>0$. Since $\textnormal{Hom}(E,\mathcal{O}_{C}(-2))=0$, we also have $\textnormal{Hom}(E_{1},\mathcal{O}_{C}(-2))=0$, hence $E_{1}\in\mathcal{T}_{\nu,-2}$ and equation 6.3.1 is also a s.e.s. in $\mathcal{B}_{\nu,-2}$. Repeating this process, we obtain $E^{\prime}\hookrightarrow L$ in $\mathcal{B}_{\nu,-2}$ and $\textnormal{Hom}(E^{\prime},\mathcal{O}_{C}(-1))=0$. Since $\mathcal{O}_{C}(-1)\in\ker Z_{V,\nu}$, we have $\sigma^{b}_{V,\nu}(E)=\sigma^{b}_{V,\nu}(E^{\prime})$.
Proposition 6.4. Let $L$ be a line bundle on $X$ with $c=0$. Let $F\hookrightarrow L$ in $\mathcal{B}_{\nu,-2}$, then there exists $0\to E\to L\to Q\to 0$ in $\mathcal{B}_{\nu,-2}$ such that $C\cdot\mathrm{ch}_{1}(H^{0}(Q))\geq-2$ and $\sigma^{b}_{V,\nu}(E)=\sigma^{b}_{V,\nu}(F)$.
@@
By the Proposition 6, there exists $0\to E\to L\to Q\to 0$ in $\mathcal{B}_{\nu,-2}$ such that $\textnormal{Hom}(E,\mathcal{O}_{C}(-1))=0$ and $\sigma^{b}_{V,\nu}(E)=\sigma^{b}_{V,\nu}(F)$. Since $H^{0}(Q)$ is a torsion sheaf, we can write $\mathrm{ch}_{1}(H^{0}(Q))=a_{0}C+\sum a_{i}C_{i}$ where $a_{i}\in\mathbb{Z}_{\geq 0}$ and $C_{i}$’s are irreducible curves in $X$ which does not equal to $C$. Then $L|_{C}\twoheadrightarrow H^{0}(Q)|_{C}$ implies that $H^{0}(Q)|_{C}\simeq\mathcal{O}_{C}$ or supported in dimension 0. If $a_{0}>0$, we have
$0\to K_{1}\to H^{0}(Q)\to\mathcal{O}_{C}\to 0,$
where $L(-C)\twoheadrightarrow K_{1}$. If nonzero, $K_{1}|_{C}\simeq\mathcal{O}_{C}(2)$ or is supported on dimension 0. If $K_{1}|_{C}\simeq\mathcal{O}_{C}(2)$, denote $\ker(K_{1}\to\mathcal{O}_{C}(2))$ by $K_{2}$ and repeat this process, we get $K_{a_{0}-1}|_{C}\simeq\mathcal{O}_{C}(2(a_{0}-1))$, and $K_{a_{0}}=0$ or supported in dimension 0. Then we have $H^{0}(Q)$ fits into a s.e.s.
$0\to M\to H^{0}(Q)\to N\to 0,$
where $M|_{C}$ is zero or supported in dimension zero, $N\in\langle\mathcal{O}_{C}(i)|i\geq 0\rangle$. Riemann-Roch computation implies that $\textnormal{Ext}^{1}(\mathcal{O}_{C}(i),\mathcal{O}_{C}(-1))\simeq\mathbb{C}^{i}$.
Claim. If $a_{0}>1$, then $\dim\textnormal{Ext}^{1}(H^{0}(Q),\mathcal{O}_{C}(-1))\geq 1$.
Indeed, it is enough to show that $\dim\textnormal{Ext}^{1}(N,\mathcal{O}_{C}(-1))\geq 1$. Hence it is enough if we have $\dim\textnormal{Ext}^{1}(N_{1},\mathcal{O}_{C}(-1))\geq 1$, where $N_{1}$ fits into the s.e.s. in $\mathrm{Coh}(X)$
$0\to\mathcal{O}_{C}(2)\to N_{1}\to\mathcal{O}_{C}\to 0.$
Let $p\in X$ such that $\textnormal{Hom}(N_{1},\mathcal{O}_{p})\simeq\mathbb{C}^{2}$. Apply $\textnormal{Hom}(\_,N_{1})$ to the s.e.s.
$0\to\mathcal{O}_{C}(-1)\to\mathcal{O}_{C}\to\mathcal{O}_{p}\to 0.$
Since $\textnormal{Ext}^{1}(\mathcal{O}_{C},N_{1})=0$, we have
$0\to\textnormal{Ext}^{1}(\mathcal{O}_{C}(-1),N_{1})\to\textnormal{Ext}^{2}(\mathcal{O}_{p},N_{1})\to\textnormal{Ext}^{2}(\mathcal{O}_{C},N_{1})\to 0.$
Hence $\dim\textnormal{Ext}^{1}(\mathcal{O}_{C}(-1),N_{1})=1$.
Denoting the image of $E\to L$ by $L_{n}(-C^{\prime})$. Then $\textnormal{Hom}(E,\mathcal{O}_{C}(-1))=0$ implies the vanishing $\textnormal{Hom}(L_{n}(-C^{\prime}),\mathcal{O}_{C}(-1))=0$ and hence $\textnormal{Ext}^{1}(H^{0}(Q),\mathcal{O}_{C}(-1))=0$. This leads to a contradiction, hence $a_{0}\leq 1$.
Corollary 6.5. Let $X$ be a smooth K3 surfaces, $C$ a (2)(-2) curve on $X$, and $L$ be a line bundle on $X$ such that $c=0$. Assume that for any irreducible curve $^{\prime\prime}\subseteq X, C′′C$$^{\prime\prime}\neq C$, we have $C^{\prime\prime}\cdot C>2$. Then $L$ is $\sigma^{b}_{V,\nu}$-semistable for all $V>0$.
@@
Assume that $E$ destabilizes $L$ at $\sigma^{b}_{V,\nu}$ for some $V$. Let $L_{n}(-C^{\prime})$ be the image of $E\to L$. By Proposition 6, we can assume that $\textnormal{Hom}(E,\mathcal{O}_{C}(-1))=0$, hence $\textnormal{Hom}(L(-C^{\prime}),\mathcal{O}_{C}(-1))=0$. This implies that $C^{\prime}\cdot C\leq 0$. On the other hand, Proposition 6 implies that $a_{0}\leq 1$. Combine these two results, for any irreducible component $C^{\prime\prime}$ of $^{\prime}, C′′C$$^{\prime\prime}\neq C$, we have $0\leq C^{\prime\prime}\cdot C\leq 2$. By assumption of the Proposition, such $C^{\prime\prime}$ does not exist, hence $C^{\prime}=C$.
Since $\phi^{b}_{V,\nu}(L_{n}(-C))\leq\phi^{b}_{V,\nu}(L)$ for any $V$, we have rank of $E$ is greater than $1$. By the discussion after Definition 6, the highest destabilizing wall for higher rank object exists. Let $F\rightarrow L$ in $\mathcal{B}_{\nu,-2}$ such that $W(F,L)=V_{h}$. Write $\mathrm{ch}_{1}(L)=m\nu+nH$, where $H\cdot\nu=0$. Consider the space of Bridgeland/weak stability conditions $\sigma^{b}_{\nu,B}$, where $B=y\nu+zH$. Note that $c=0$ implies that $H\cdot C=0$ and $B\cdot C=0$ for any $y,z$. Hence $\sigma^{b}_{\nu,B}$ has the heart $\mathcal{B}^{y}_{\nu,-2}$. Furthermore, since $W(F,L)=V_{h}$, $F$ must be torsion free, and hence a suboject of $L$ in $\mathrm{Coh}^{y}$. Indeed, if $F$ is not torsion free, denoting the torsion subsheaf of $F$ by $F_{0}$, then $F/F_{0}$ is a subobject of $L$ in $\mathcal{B}^{y}_{\nu,-2}$, and $W(F/F_{0},L)_{z}$ is higher than $W(F,L)_{z}$, contradicting $W(F,L)$ is the highest wall at $(y,z)$. The same argument in [2] implies that $W(F,L)$ is lower than $W(L_{n}(-C),L)$ for $|z|\gg 0$, which contradicts $\phi^{b}_{V,\nu,B}(L_{n}(-C))<\phi^{b}_{V,\nu,B}(L)$ for any $y,z$. Hence such $F$ does not exists.
The K3 surfaces satisfies the assumption in the Corollary 6 exists.
@@
Example 6.6. Consider the case Theorem 2(v) in [11]. Let $\rho(X)=2$. Then the effective cone of $X$ is generated by two smooth rational curves of self intersection 2-2. Denoting the two curves by $C_{1}$ and $C_{2}$ and $C_{1}\cdot C_{2}=q$, then $q>2$. Since otherwise, the first Chern character of any line bundle has non-positive self-intersection, which implies that there is no ample line bundle on $X$, a contradiction. In this case, take $y$ to be an even positive integer, and $\nu=\frac{q}{2}yC_{1}+yC_{2}$, then $\nu$ satisfies $\nu\cdot C_{1}=0$, $\nu\cdot C_{2}>0$ and $\nu^{2}>0$.
Example 6.7. Again consider the case Theorem 2(v) in [11]. Let $\rho(X)=3$. Assume that the effective cone of $X$ is generated by three curves $C_{1}$, $C_{2}$ and $C_{3}$. Then Corollary 2.9 (i) in [13] implies that there exists K3 surface such that the pairwise intersection between $C_{i}$’s are greater than $2$.